40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 524     Entry time: Fri Jun 6 16:10:51 2008
Author: steve 
Type: Update 
Category: PSL 
Subject: HEPA filters are running at 100% 
The psl HEPA filters were turned up to run at 100% to accommodate beam trap work on Tuesday, June 3, 2008
ELOG V3.1.3-